OpenStudy (anonymous):

f(x)=x2-3x+3 g(x)x-1/4 x=5 Find f(g(x)

OpenStudy (jdoe0001):

\(\large g(x)x-1/4 x=5 \quad ?\)

OpenStudy (anonymous):

x-1/4 ....x = 5

OpenStudy (jdoe0001):

hmmm... can you post a quick screenshot of it?

OpenStudy (anonymous):

They wont let me >/ its x-1 divided by 4

OpenStudy (jdoe0001):

they won't let... you? who? press PrintScr button, next to the F12 key then go to Paint and Paste there, save as gif or jpeg and post here

OpenStudy (rsadhvika):

\(f(x)=x^2-3x+3 \\ g(x) = \frac{x-1}{4} \\ x = 5 \)

OpenStudy (rsadhvika):

like above ?

OpenStudy (rsadhvika):

and you want to find \(f(g(x)) \) Is that right ?

OpenStudy (anonymous):

Yes correct !

OpenStudy (rsadhvika):

good, basically they want you find \(f(g(5))\) so, first find \(g(5)\) \(g(5) = \frac{5-1}{4} = ?\)

OpenStudy (anonymous):

1!

OpenStudy (rsadhvika):

Yes, \(g(5)=1\) so, \(f(g(5))=f(1)=1^2−3(1)+3=?\)

OpenStudy (anonymous):

-5?

OpenStudy (rsadhvika):

try again :)

OpenStudy (rsadhvika):

\(f(g(5))=f(1)=1^2−3(1)+3= 1 -3 +3 = ?\)

OpenStudy (anonymous):

1!

OpenStudy (rsadhvika):

yup !

OpenStudy (anonymous):

Thank youuu :)

OpenStudy (anonymous):

can u help with another one please @rsadhvika